# Generalization bounds for mixing processes via delayed online-to-PAC conversions

### v1

@@ Abstract
We study the generalization error of statistical learning algorithms in a non-i.i.d. setting, where the training data is sampled from a stationary mixing process. We develop an analytic framework for this scenario based on a reduction to online learning with delayed feedback. In particular, we show that the existence of an online learning algorithm with bounded regret (against a fixed statistical learning algorithm in a specially constructed game of online learning with delayed feedback) implies low generalization error of said statistical learning method even if the data sequence is sampled from a mixing time series. The rates demonstrate a trade-off between the amount of delay in the online learning game and the degree of dependence between consecutive data points, with near-optimal rates recovered in a number of well-studied settings when the delay is tuned appropriately as a function of the mixing time of the process.
@@ 1 Introduction
In machine learning, generalization means the ability of a model to infer patterns from a dataset of training examples and apply them to analyze previously unseen data (Shalev-Shwartz and Ben-David,, 2014). The gap in accuracy between the model’s predictions on new data and those on the training set is usually referred to as generalization error. Providing upper bounds on this quantity is a central goal in statistical learning theory. Classically, bounds based on notions of complexity of the model’s hypothesis space (e.g., its VC dimension or Rademacher complexity) were used to provide uniform worst-case guarantees (see Bousquet et al.,, 2004; Vapnik,, 2013; Shalev-Shwartz and Ben-David,, 2014). However, results of this kind are often too loose to be applied to the most common machine learning over-parameterised models, such as deep neural networks (Zhang et al.,, 2021). As a consequence, several approaches have been proposed to obtain algorithm-dependent generalization bounds, which can adapt to the problem and be much tighter in practice than their uniform counterparts. Often, the underlying idea is that if the algorithm’s output does not have a too strong dependence on the specific input dataset used for the training, then the model should not be prone to overfitting, and thus generalize well. Examples of results that build onto these ideas are stability bounds, information-theoretic bounds, and PAC-Bayesian bounds (see, e.g., Bousquet and Elisseeff,, 2002; Russo and Zou,, 2020; Hellström et al.,, 2023; Alquier,, 2024).
@@
Most results in the literature focus on the i.i.d. setting, where the training dataset is made of independent draws from some underlying data distribution. However, for several applications, this assumption is far from realistic. For instance, it excludes the case where observations received by the learner have some inherent temporal dependence, as it is the case for stock prices, daily energy consumption, or sensor data from physical environments (Ariyo et al.,, 2014; Takeda et al.,, 2016). This calls for the development of theory for addressing non-i.i.d. data. A common approach in the extant literature is to consider a class of non-i.i.d. data-generating processes usually referred to as stationary $\beta$-mixing or $\varphi$-mixing processes. This assumption, together with a “blocking” trick introduced by Yu, (1994), has led to a few results in the literature: Meir, (2000), Mohri and Rostamizadeh, (2008), Shalizi and Kontorovich, (2013), and Wolfer and Kontorovich, (2019) provided uniform worst-case generalization bounds, Steinwart and Christmann, (2009) and Agarwal and Duchi, (2012) discussed excess risk bounds (comparing the algorithm’s output with the best possible hypothesis), while Mohri and Rostamizadeh, (2010) gave bounds based on a stability analysis (in the sense of Bousquet and Elisseeff,, 2002).
@@
In the present paper, we propose a new framework for proving generalization bounds for the non-i.i.d. setting, which take a form that is similar in spirit to PAC-Bayesian bounds (Guedj,, 2019; Alquier,, 2024): high-probability upper bounds on the expected generalization error of randomized learning algorithms. We achieve this by combining the blocking argument by Yu, (1994) to manage the concentration of sums of correlated random variables, with the recent online-to-PAC conversion technique recently proposed by Lugosi and Neu, (2023). Using their framework we show a new way to obtain generalization bounds for stationary dependent processes that satisfy a certain “short-memory” property, intuitively meaning that data points that are closer in time are more heavily dependent on each other. Our assumption slightly differs from $\beta$-mixing in the sense that we only need it to hold for a specific class of bounded loss functions (as opposed to requiring strong conditions phrased in terms of total variation distance). Among other results, this allows us to prove PAC-Bayesian generalization bounds for mixing processes. This complements previous work on such bounds that have only considered mild relaxations of the i.i.d. condition such as assuming that the data has a martingale structure (see, e.g., Seldin et al.,, 2012; Chugg et al.,, 2023; Haddouche and Guedj,, 2023). Notable exceptions are the works of Alquier and Wintenberger, (2012), Alquier et al., (2013), and Eringis et al., (2022, 2024), who provided generalization bounds for a sequential prediction setting where both the data-generating process and the hypothesis class used for prediction are stable dynamical systems. Their results are proved under some very specific conditions on these systems, and their guarantees involve unspecified problem-dependent constants that may be large. In contrast, our bounds hold under general, simple-to-verify conditions and feature explicit constants. The rest of the paper is organized as follows. In Section 2 we properly define the generalization error of a statistical learning algorithm for both i.i.d. and non-i.i.d. cases, and state our main assumption on the data dependence. Our main contribution lies in Section 3, where after recalling the results of Lugosi and Neu, (2023) for the i.i.d. setting we show how to adapt their technique to stationary mixing processes. In Section 4 we provide some concrete results of the bounds we can obtain through the resulting online-to-PAC conversion methodology. Finally in Section 5 we extend our results to the setting where the hypothesis class itself may consist of dynamical systems.
@@ 2 Preliminaries
The classical statistical learning framework usually considers a dataset $S_{n}=(Z_{1},...,Z_{n})$, made of $n$ i.i.d. elements drawn from a distribution $\mu$ over a measurable instance space $\mathcal{Z}$. Furthermore, we are given a measurable class $\mathcal{W}$ of hypotheses and a loss function $\ell:\mathcal{W}\times\mathcal{Z}\to{}_{+}$, with $\ell(w,z)$ measuring the quality of the hypothesis $w\in\mathcal{W}$ on the data instance $z\in\mathcal{Z}$. For any given hypothesis $w\in\mathcal{W}$, two key objects of interest are the training error ^$\widehat{\mathcal{L}}(w,S_{n})=\frac{1}{n}\sum_{i=1}^{n}\ell(w,Z_{i})$ and the test error $\mathcal{L}(w)=\mathbb{E}_{Z^{\prime}\sim\mu}[\ell(w,Z^{\prime})]$, where the random element $Z^{\prime}$ has the same distribution as $Z_{i}$ and is independent of $S_{n}$.
A learning algorithm $\mathcal{A}:\mathcal{Z}^{n}\to\mathcal{W}$ maps the training sample to a hypothesis in $\mathcal{W}$. More generally, we will focus on randomized learning algorithms, returning a probability distribution $P_{W_{n}|S_{n}}\in\Delta_{\mathcal{W}}$ over $\mathcal{W}$, conditionally on $S_{n}$ (with deterministic algorithms recovered as special cases, whose the outputs are Dirac distributions). The ultimate goal of the learner is to minimize the test error. Yet, this quantity cannot be computed without knowledge of the data generating distribution $\mu$. In practice, one typically relies on the training error in order to gauge the quality of the algorithm. For an algorithm $\mathcal{A}:S_{n}\mapsto P_{W_{n}|S_{n}}$, we define the generalization error as the expected gap between training and test error:
@@
The expectation in the above expression integrates over the randomness in the output of the algorithm $W_{n}\sim P_{W_{n}|S_{n}}$, conditionally on the sample $S_{n}$. We remark that the test error is generally not equal to the mean of the training error, due to the dependence of $W_{n}$ on the training data.
@@
We extend the previous setting by considering the case where the data have an intrinsic temporally ordered structure, and come in the form of a stationary process $(Z_{t})_{t\in\mathbb{N}^{*}}\sim\nu$. Formally, we assume that the joint marginal distribution of any block $(Z_{t},Z_{t-1},\dots,Z_{t-i})$ is the same as the distribution of $(Z_{t+j},Z_{t+j-1},\dots,Z_{t+j-i})$ for any $t$, $i$ and $j$, but the data points are not necessarily independent of each other. In particular, the marginal distribution of $Z_{t}$ is constant and is denoted by $\mu$. Thus, it is natural to continue to use the definition of the test loss and generalization error given above, although with the understanding that $\mu$ now refers to the marginal distribution of an independent copy of $Z_{1}$, a sample point from a stationary non-i.i.d. process. We will use $\mathcal{L}(w)=\mathbb{E}_{Z^{\prime}\sim\mu}[\ell(w,Z^{\prime})]$ to denote the test loss defined in this sense. We remark here that other notions of the test loss may also be considered, and the framework that we propose can be extended to most natural definitions with little work (but potentially large notational overhead). In Section 5, we provide such an extension for a more general setting where the hypotheses themselves are allowed to have memory and the process may not be as strongly stationary as our assumption above requires.
@@
In order to obtain generalization bounds we need to have some control on how strong the dependencies between different datapoints are allowed to be. To this regard, we consider the following assumption.
Assumption 1.
@@
There exists a non-increasing sequence $(\phi_{d})_{d\in\mathbb{N}^{*}}$ of non-negative real numbers such that, for all $w\in\mathcal{W}$ and all $t,d\in\mathbb{N}^{*}$, the following holds:
$\mathbb{E}\left[\mathcal{L}(w)-\ell(w,Z_{t})\Big{|}\mathcal{F}_{t-d}\right]\leq\phi_{d}\,.$
@@
The intuition behind this assumption is that the loss associated with the observations $Z_{t}$ becomes almost independent of the past after $d$ steps, enabling us to treat each sequence of the form $(Z_{t},Z_{t+d},\dots,Z_{t+(n-t)d})$ as an approximately i.i.d. sequence. Note that this assumption differs from the usual $\beta$-mixing assumption which requires the distribution of $Z_{t}|\mathcal{F}_{t-d}$ to be close to the marginal distribution $\mu$ for all $t$, in terms of total variation distance. Our assumption is somewhat weaker in the sense that it only requires the expected losses under these distributions to be close, and only a one-sided inequality is required. It is easy to verify that our assumption is satisfied if the process is $\beta$-mixing in the usual sense and the losses are bounded in $[0,1]$.
@@ 3 Proving generalization bounds via online learning
This section introduces our main contribution: a framework for proving generalization bounds for statistical learning on non-i.i.d. data via a reduction to online learning. The framework of online learning focuses on algorithms that aim to improve performance (measured in terms of a given cost function) incrementally as new information becomes available, often without any underlying assumption on how data are generated. The online learner’s performance is typically measured in terms of its regret, defined as the the difference between the cumulative cost of the online learner and that of a fixed comparator. We refer to the monographs Cesa-Bianchi and Lugosi,, 2006 and Orabona,, 2019 for comprehensive overviews on online learning and regret analysis. Recently, Lugosi and Neu, (2023) established a connection between upper bounds on the regret and generalization bounds, showing that the existence of a strategy with a bounded regret in a specially designed online game translates into a generalization bound, via a technique dubbed online-to-PAC conversion. Their focus is on the i.i.d. setting, where the training dataset is made of independent draws. Here, we show that this framework can naturally be extended beyond the i.i.d. assumption.
In what follows, we briefly review the setup of Lugosi and Neu, (2023) in Section 3.1 and then describe our new extension of their model to the non-i.i.d. case in Section 3.2. In particular, we prove a high-probability bound for the generalization error of any statistical learning algorithm learnt with a stationary mixing process verifying Assumption 1.
@@
The main idea of Lugosi and Neu, (2023) is to introduce an online learning game (associated with the statistical learning setup) called the generalization game, where the following steps are repeated for a sequence of rounds $t=1,2,\dots,n$:
  • The online learner picks a distribution $P_{t}\in\Delta_{\mathcal{W}}$,
  • the adversary selects the cost function $c_{t}:w\mapsto\ell(w,Z_{t})-\mathcal{L}(w)$,
  • the online learner incurs the cost $\langle P_{t},c_{t}\rangle=\mathbb{E}_{W\sim P_{t}}[c_{t}(W)]$,
@@
Theorem 1 (Theorem 1 in Lugosi and Neu,, 2023; see appendix A.1).
With the notation introduced above, the generalization error of algorithm $\mathcal{A}$ on the training set $S_{n}$ can be written as
@@
As a concrete application of the above, the following generalization bound is obtained when picking the classic exponential weighted average (EWA) algorithm (Vovk,, 1990; Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997) as online strategy, and plugging its regret bound into (1).
Theorem 2 (Corollary $6$ in Lugosi and Neu,, 2023).
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathcal{D}_{\mathrm{KL}}\left(P_{W_{n}|S_{n}}\middle\|P_{1}\right)}{\eta n}+\frac{\eta}{2}+\sqrt{\frac{2\log\left(\frac{1}{\delta}\right)}{n}}\,.$
Proof.
We can bound each term of (1) separately. A data-dependent bound for the regret term is obtained via a direct application of the regret analysis of EWA which brings the term $\frac{\mathcal{D}_{\mathrm{KL}}\left(P_{W_{n}|S_{n}}\middle\|P_{1}\right)}{\eta n}+\frac{\eta}{2}$ (see Appendix B.1). The term $\sqrt{\frac{2\log\left(\frac{1}{\delta}\right)}{n}}$ results from bounding the martingale $M_{n}$ via an application of Hoeffding–Azuma inequality. ∎
Note that the first term in the above bound is data-dependent due to the presence of $P_{W_{n}|S_{n}}$, and thus optimizing it requires a data-dependent choice of $\eta$, which is not allowed by Theorem 2. However, via a standard covering argument it is possible to get a bound in the form
$\mathrm{Gen}(\mathcal{A},S_{n})=\mathcal{O}\left(\sqrt{\frac{\mathcal{D}_{\mathrm{KL}}\left(P_{W_{n}|S_{n}}\middle\|P_{1}\right)}{n}}+\sqrt{\frac{1}{n}\log\left(\frac{\log n}{\delta}\right)}\right).$
For the details, we refer to the proof of Corollary 5 of Lugosi and Neu, (2023), which recovers a classical PAC-Bayes bound of McAllester, (1998).
@@
We will now drop the i.i.d. assumption for the data, and instead consider non-i.i.d. sequences satisfying Assumption 1. For this setting we define the following variant of the generalization game.
Definition 1 (Generalization game with delay).
@@
  • The online learner picks a distribution $P_{t}\in\Delta_{\mathcal{W}}$,
  • the adversary selects the cost function $c_{t}:w\mapsto\ell(w,Z_{t})-\mathcal{L}(w)$,
  • the online learner incurs the cost $\langle P_{t},c_{t}\rangle=\mathbb{E}_{W\sim P_{t}}[c_{t}(W)]$,
  • if $t\geq d$, then $Z_{t-d+1}$ is revealed to the learner.
As before, the last step effectively reveals the entire cost function $c_{t-d+1}$ to the online learner. The main difference between our version of the generalization game and the standard one of Lugosi and Neu, (2023) is the introduction of a delay on the online learning algorithm’s decisions. Specifically, we will force the online learner to only take information into account up to time $t-d$ when picking their action $P_{t}$. Clearly, setting $d=1$ recovers the original version of the generalization game.
It is easy to see that the regret decomposition of Theorem 1 still remains valid in the current setting. The purpose of introducing the delay is to be able to make sure that the term $M_{n}=-\frac{1}{n}\sum_{t=1}^{n}\left\langle P_{t},c_{t}\right\rangle$ is small. The lemma below states that the increments of $M_{n}$ behave similarly to a martingale-difference sequence, thanks to the introduction of the delay.
Lemma 1.
Fix $d\in[\![1,n]\!]$. Under assumption 1, it holds for all $t\in[\![1,n]\!]$:
@@
where $P_{t}$ and $c_{t}$ are defined as in 1.
Proof.
Since $P_{t}$ is $\mathcal{F}_{t-d}$-measurable we have $\mathbb{E}[\langle-P_{t},c_{t}\rangle|\mathcal{F}_{t-d}]=\langle P_{t},\mathbb{E}[-c_{t}|\mathcal{F}_{t-d}]\rangle\leq\phi_{d}$, where the last step uses Assumption 1. ∎
Thus, by following the decomposition of Theorem 1, we are left with the problem of bounding the regret of the delayed online learning algorithm against $P_{W_{n}|S_{n}}$, denoted as $\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})=\sum_{t=1}^{n}\left\langle P_{t}-P_{W_{n}|S_{n}},c_{t}\right\rangle$. The following proposition states a simple and clean bound that one can immediately derive from these insights.
Proposition 1 (Bound in expectation).
@@
Proof.
By Theorem 1, it holds that $\mathbb{E}[\mathrm{Gen}(\mathcal{A},S_{n})]=\frac{\mathbb{E}[\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})]}{n}+\mathbb{E}[M_{n}]$, where the regret is for a strategy $P_{t}$ in the delayed generalization game. Hence, by Lemma 1
@@
which proves the claim. ∎

### v3

@@ Abstract
We study the generalization error of statistical learning algorithms in a non i.i.d. setting, where the training data is sampled from a stationary mixing process. We develop an analytic framework for this scenario based on a reduction to online learning with delayed feedback. In particular, we show that the existence of an online learning algorithm with bounded regret (against a fixed statistical learning algorithm in a specially constructed game of online learning with delayed feedback) implies low generalization error of said statistical learning method even if the data sequence is sampled from a mixing time series. The rates demonstrate a trade-off between the amount of delay in the online learning game and the degree of dependence between consecutive data points, with near-optimal rates recovered in a number of well-studied settings when the delay is tuned appropriately as a function of the mixing time of the process.
@@ 1 Introduction
In machine learning, generalization means the ability of a model to infer patterns from a dataset of training examples and apply them to analyze previously unseen data (Shalev-Shwartz and Ben-David, 2014). The gap in accuracy between the model’s predictions on new data and those on the training set is usually referred to as generalization error. Providing upper bounds on this quantity is a central goal in statistical learning theory. Classically, bounds based on notions of complexity of the model’s hypothesis space (e.g., its VC dimension or Rademacher complexity) were used to provide uniform worst-case guarantees (see Bousquet et al., 2004; Vapnik, 2013; Shalev-Shwartz and Ben-David, 2014). However, these results are often too loose to be applied to over-parameterised models such as deep neural networks (Zhang et al., 2021). As a consequence, several approaches have been proposed to obtain algorithm-dependent generalization bounds, which can adapt to the problem and be much tighter in practice than their uniform counterparts. Often, the underlying idea is that if the algorithm’s output does not have a too strong dependence on the specific input dataset used for the training, then the model should not be prone to overfitting, and thus generalize well. Examples of results that build onto these ideas are stability bounds, information-theoretic bounds, and PAC-Bayesian bounds (see, e.g., Bousquet and Elisseeff, 2002; Russo and Zou, 2020; Hellström et al., 2023; Alquier, 2024).
@@
Most results in the literature focus on the i.i.d. setting, where the training dataset is made of independent draws from some underlying data distribution. However, for several applications, this assumption is far from realistic. For instance, it excludes the case where observations received by the learner have some inherent temporal dependence, as it is the case for stock prices, daily energy consumption, or sensor data from physical environments (Ariyo et al., 2014; Takeda et al., 2016). This calls for the development of theory for addressing non-i.i.d. data. A common approach in the extant literature is to consider a class of non-i.i.d. data-generating processes usually referred to as stationary $\beta$-mixing or $\varphi$-mixing processes. This assumption, together with a “blocking” trick introduced by Yu (1994), has led to a few results in the literature: Meir (2000), Mohri and Rostamizadeh (2008), Shalizi and Kontorovich (2013), and Wolfer and Kontorovich (2019) provided uniform worst-case generalization bounds, Steinwart and Christmann (2009) and Agarwal and Duchi (2012) discussed excess risk bounds (comparing the algorithm’s output with the best possible hypothesis), while Mohri and Rostamizadeh (2010) gave bounds based on a stability analysis (in the sense of Bousquet and Elisseeff, 2002).
@@
In the present paper, we propose a new framework for proving generalization bounds for the non-i.i.d. setting, which take a form that is similar in spirit to PAC-Bayesian bounds (Guedj, 2019; Alquier, 2024): high-probability upper bounds on the expected generalization error of randomized learning algorithms. We achieve this by combining the blocking argument by Yu (1994) to manage the concentration of sums of correlated random variables, with the recent online-to-PAC conversion technique recently proposed by Lugosi and Neu (2023). Using their framework we show a new way to obtain generalization bounds for stationary dependent processes that satisfy a certain “short-memory” property, intuitively meaning that data points that are closer in time are more heavily dependent on each other. Our assumption slightly relaxes the commonly considered $\beta$-mixing condition in that we only need it to hold for a specific class of bounded loss functions (as opposed to requiring stronger conditions phrased in terms of total variation distance). Among other results, this allows us to prove PAC-Bayesian generalization bounds for mixing processes. This complements previous work on such bounds that have largely considered mild relaxations of the i.i.d. condition such as assuming that the data has a martingale structure (e.g., Seldin et al., 2012; Chugg et al., 2023; Haddouche and Guedj, 2023). A notable exception is the work of Ralaivola et al. (2010), who considered data sets whose dependence structure is described via graphs, and proved generalization bounds that scale with various graph properties describing the overall strength of dependencies. One of the applications of their results is a generalization bound for $\beta$-mixing processes, comparable in nature to some of our results. We return to discussing the similarities and differences in Section 6. Further relevant works are those of Alquier and Wintenberger (2012), Alquier et al. (2013), and Eringis et al. (2022, 2024), who provided generalization bounds for a sequential prediction setting where both the data-generating process and the hypothesis class used for prediction are stable dynamical systems. Their results are proved under some very specific conditions on these systems, and their guarantees involve unspecified problem-dependent constants that may be large. In contrast, our bounds hold under general, simple-to-verify conditions and feature explicit constants.
The rest of the paper is organized as follows. In Section 2 we properly define the generalization error of a statistical learning algorithm for both i.i.d. and non-i.i.d. cases, and state our main assumption on the data dependence. Our main contribution lies in Section 3, where after recalling the results of Lugosi and Neu (2023) for the i.i.d. setting we show how to adapt their technique to stationary mixing processes. In Section 4 we provide some concrete results of the bounds we can obtain through the resulting online-to-PAC conversion methodology. Finally in Section 5 we extend our results to the setting where the hypothesis class itself may consist of dynamical systems.
@@ 2 Preliminaries
The classical statistical learning framework usually considers a dataset $S_{n}=(Z_{1},...,Z_{n})$, made of $n$ i.i.d. elements drawn from a distribution $\mu$ over a measurable instance space $\mathcal{Z}$. Often, one can think of each $Z_{i}$ as a feature-label pair $(X_{i},Y_{i})$. Furthermore, we are given a measurable class $\mathcal{W}$ of hypotheses and a loss function $\ell:\mathcal{W}\times\mathcal{Z}\to\mathbb{R}_{+}$, with $\ell(w,z)$ measuring the quality of the hypothesis $w\in\mathcal{W}$ on the data instance $z\in\mathcal{Z}$. For any given hypothesis $w\in\mathcal{W}$, two key objects of interest are the training error ^$\widehat{\mathcal{L}}(w,S_{n})=\frac{1}{n}\sum_{i=1}^{n}\ell(w,Z_{i})$ and the test error $\mathcal{L}(w)=\mathbb{E}_{Z^{\prime}\sim\mu}[\ell(w,Z^{\prime})]$, where the random element $Z^{\prime}$ has the same distribution as $Z_{i}$ and is independent of $S_{n}$. A learning algorithm $\mathcal{A}:\mathcal{Z}^{n}\to\mathcal{W}$ maps the training sample to an hypothesis in $\mathcal{W}$. More generally, we will focus on randomized learning algorithms, returning a probability distribution $P_{W_{n}|S_{n}}\in\Delta_{\mathcal{W}}$ over $\mathcal{W}$, conditionally on $S_{n}$ (deterministic algorithms can be recovered as special cases, whose the outputs are Dirac distributions). The ultimate goal of the learner is to minimize the test error. Yet, this quantity cannot be computed without knowledge of the data generating distribution $\mu$. In practice, one typically relies on the training error in order to gauge the quality of the algorithm. For an algorithm $\mathcal{A}:S_{n}\mapsto P_{W_{n}|S_{n}}$, we define the generalization error as the expected gap between training and test error:
@@
The expectation in the above expression integrates over the randomness in the output of the algorithm $W_{n}\sim P_{W_{n}|S_{n}}$, conditionally on the sample $S_{n}$. We stress that the test error is generally not equal to the mean of the training error, due to the dependence of $W_{n}$ on the training data, which is precisely the challenge that necessitates studying conditions under which the generalization error is small.
@@
We extend the previous setting by considering the case where the data have an intrinsic temporally ordered structure, and come in the form of a stationary process $(Z_{t})_{t\in\mathbb{N}^{*}}\sim\nu$. Formally, we assume that the joint marginal distribution of any block $(Z_{t},Z_{t-1},\dots,Z_{t-i})$ is the same as the distribution of $(Z_{t+j},Z_{t+j-1},\dots,Z_{t+j-i})$ for any $t$, $i$ and $j$, but the data points are not necessarily independent of each other. In particular, the marginal distribution of $Z_{t}$ is constant and is denoted by $\mu$. In such a setting, it is natural to continue to use the definition of the test loss and generalization error given above, although with the understanding that $\mu$ now refers to the marginal distribution of an independent copy of $Z_{1}$, a sample point from a stationary non-i.i.d. process. We remark here that other notions of the test loss may also be considered, and the framework that we propose can be extended to most natural definitions with little work (but potentially large notational overhead). In Section 5, we provide such an extension for a more general setting where the hypotheses themselves are allowed to have memory and the process may not be as strongly stationary as our assumption above requires.
@@
To obtain generalization results we need some control on how strong the dependencies between different datapoints are allowed to be. We hence consider the following assumption.
Assumption 1
@@
There exists a non-increasing sequence $(\phi_{d})_{d\in\mathbb{N}^{*}}$ of non-negative real numbers such that, for all $w\in\mathcal{W}$ and all $t\in\mathbb{N}^{*}$:
$\mathbb{E}\left[\mathcal{L}(w)-\ell(w,Z_{t})\Big{|}\mathcal{F}_{t-d}\right]\leq\phi_{d}\,,$
where $\mathcal{L}(w)=\mathbb{E}_{Z^{\prime}\sim\mu}[\ell(w,Z^{\prime})]$, with $Z^{\prime}$ being independent on the process $(Z_{t})_{t\in\mathbb{N}^{*}}$ and having as distribution the stationary marginal $\mu$ of the $Z_{t}$.
@@
The intuition behind this assumption is that the loss associated with the observations $Z_{t}$ becomes almost independent of the past after $d$ steps, enabling us to treat each sequence of the form $(Z_{t},Z_{t+d},\dots,Z_{t+(n-t)d})$ as an approximately i.i.d. sequence. Note that this assumption differs from the usual $\beta$-mixing assumption that requires the distribution of $Z_{t}|\mathcal{F}_{t-d}$ to be close (in total variation) to the marginal distribution $\mu$ for all $t$. Our assumption is somewhat weaker in the sense that it only requires the expected losses under these distributions to be close, and only a one-sided inequality is required. It is easy to verify that our assumption is satisfied if the process is $\beta$-mixing in the usual sense and the losses are bounded in $[0,1]$.
It is worth noticing that while our assumption can look like a small “cosmetic” improvement over the standard $\beta$-mixing, it can actually be much weaker. For a concrete example, consider let $\ell(w,Z_{t})=\ell(w,Z^{\prime}_{t})+\varepsilon_{t}$, where the $Z_{t}^{\prime}$ are part of an i.i.d. sequence, and $\varepsilon_{t}$ is sampled from a bounded $\beta$-mixing process, with $\alpha$ a small constant. Now, for any $w$, $\ell(w,Z_{t})$ is clearly a $\beta$-mixing process (inheriting the properties of $\varepsilon_{t}$), independently of the choice of $\alpha$. In contrast, the mixing rate as defined in our assumption improves linearly with $\alpha$, and vanishes as this parameter approaches zero. As a matter of fact, $\beta$-mixing conditions are very strict in that they require mixing in terms of the entire distribution of the loss, and ignores the “scale” at which non-stationarity impacts the outcomes.
@@ 3 Proving generalization bounds via online learning
This section introduces our main contribution: a framework for proving generalization bounds for statistical learning on non-i.i.d. data via a reduction to online learning. The framework of online learning focuses on algorithms that aim to improve performance (measured in terms of a given cost function) incrementally as new information becomes available, often without any underlying assumption on how data are generated. The online learner’s performance is typically measured in terms of its regret, defined as the the difference between the cumulative cost of the online learner and that of a fixed comparator. We refer to the monographs Cesa-Bianchi and Lugosi (2006) and Orabona (2019) for comprehensive overviews on online learning and regret analysis. Recently, Lugosi and Neu (2023) established a connection between upper bounds on the regret and generalization bounds, showing that the existence of a strategy with a bounded regret in a specially designed online game translates into a generalization bound, via a technique dubbed online-to-PAC conversion. Their focus is on the i.i.d. setting, where the training dataset is made of independent draws. Here, we show that this framework can naturally be extended beyond the i.i.d. assumption.
In what follows, we briefly review the setup of Lugosi and Neu (2023) in Section 3.1 and then describe our new extension of their model to the non-i.i.d. case in Section 3.2. In particular, we prove a high-probability bound for the generalization error of any statistical learning algorithm learnt with a stationary mixing process verifying Assumption 1.
@@
The main idea of Lugosi and Neu (2023) is to introduce an online learning generalization game, where the following steps are repeated for a sequence of rounds $t=1,2,\dots,n$:
  • the online learner picks a distribution $P_{t}\in\Delta_{\mathcal{W}}$;
  • the adversary selects the cost function $c_{t}:w\mapsto\ell(w,Z_{t})-\mathcal{L}(w)$;
  • the online learner incurs the cost $\langle P_{t},c_{t}\rangle=\mathbb{E}_{W\sim P_{t}}[c_{t}(W)]$;
@@
Theorem 3.1 (Theorem 1 in Lugosi and Neu, 2023; see appendix A.1).
With the notation introduced above,
@@
As a concrete application of the above, the following generalization bound is obtained when picking the classic exponential weighted average (EWA) algorithm (Vovk, 1990; Littlestone and Warmuth, 1994; Freund and Schapire, 1997) as our online-learning strategy, and plugging its regret bound into Equation (1).
Theorem 3.2 (Corollary $6$ in Lugosi and Neu, 2023).
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathcal{D}_{KL}(P_{W_{n}|S_{n}}||P_{1})}{\eta n}+\frac{\eta}{2}+\sqrt{\frac{2\log\left(\frac{1}{\delta}\right)}{n}}\,.$
Proof 3.3.
We can bound each term of (1) separately. A data-dependent bound for the regret term is obtained via a direct application of the regret analysis of EWA which brings the term $\frac{\mathcal{D}_{KL}(P_{W_{n}|S_{n}}||P_{1})}{\eta n}+\frac{\eta}{2}$ (see Appendix B.1). The term $\sqrt{\frac{2\log\left(\frac{1}{\delta}\right)}{n}}$ results from bounding the martingale $M_{n}$ via an application of Hoeffding–Azuma inequality.
Note that the first term in the bound is data-dependent due to the presence of $P_{W_{n}|S_{n}}$, and thus optimizing it requires a data-dependent choice of $\eta$, which is not allowed by Theorem 3.2. However, via a union bound argument it is possible to get a bound of the form
$\mathrm{Gen}(\mathcal{A},S_{n})=\mathcal{O}\left(\sqrt{\frac{\mathcal{D}_{KL}(P_{W_{n}|S_{n}}||P_{1})}{n}}+\sqrt{\frac{1}{n}\log\left(\frac{\log n}{\delta}\right)}\right),$
For the details, we refer to the proof of Corollary 5 of Lugosi and Neu (2023), which recovers a classical PAC-Bayes bound of McAllester (1998).
@@
We will now drop the i.i.d. condition, and instead consider non-i.i.d. data sequences satisfying Assumption 1. For this setting we define the following variant of the generalization game.
Definition 3.4 (Generalization game with delay).
@@
  • the online learner picks a distribution $P_{t}\in\Delta_{\mathcal{W}}$;
  • the adversary selects the cost function $c_{t}:w\mapsto\ell(w,Z_{t})-\mathcal{L}(w)$;
  • the online learner incurs the cost $\langle P_{t},c_{t}\rangle=\mathbb{E}_{W\sim P_{t}}[c_{t}(W)]$;
  • if $t\geq d$, $Z_{t-d+1}$ is revealed to the learner.
As before, the last step effectively reveals the entire cost function $c_{t-d+1}$ to the online learner. The main difference between our version of the generalization game and the standard one of Lugosi and Neu (2023) is the introduction of a delay on the online learning algorithm’s decisions. Specifically, we will force the online learner to only take information into account up to time $t-d$ when picking their action $P_{t}$. Clearly, setting $d=1$ recovers the original version of the generalization game.
It is easy to see that the regret decomposition of Theorem 3.1 still remains valid in the current setting. The purpose of introducing the delay is to be able to make sure that the term $M_{n}=-\frac{1}{n}\sum_{t=1}^{n}\left\langle P_{t},c_{t}\right\rangle$ is small. The lemma below states that the increments of $M_{n}$ behave similarly to a martingale-difference sequence, thanks to the introduction of the delay.
Lemma 3.5.
Fix $d\in[\![1,n]\!]$. Under Assumption 1, defining $P_{t}$ and $c_{t}$ as in 3.4, for all $t\in[\![1,n]\!]$
@@
Proof 3.6.
Since $P_{t}$ is $\mathcal{F}_{t-d}$-measurable we have $\mathbb{E}[\langle-P_{t},c_{t}\rangle|\mathcal{F}_{t-d}]=\langle P_{t},\mathbb{E}[-c_{t}|\mathcal{F}_{t-d}]\rangle\leq\phi_{d}$, where the last step uses Assumption 1.
Thus, by following the decomposition of Theorem 3.1, we are left with the problem of bounding the regret of the delayed online learning algorithm against $P_{W_{n}|S_{n}}$, denoted as $\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})=\sum_{t=1}^{n}\left\langle P_{t}-P_{W_{n}|S_{n}},c_{t}\right\rangle$. The following proposition states a simple and clean bound that one can immediately derive from these insights.
Proposition 3.7 (Bound in expectation).
@@
Proof 3.8.
By Theorem 3.1, it holds that $\mathbb{E}[\mathrm{Gen}(\mathcal{A},S_{n})]=\frac{\mathbb{E}[\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})]}{n}+\mathbb{E}[M_{n}]$, where the regret is for a strategy $P_{t}$ in the delayed generalization game. Hence, by Lemma 3.5
@@
which proves the claim.